Two blocks connected by a string are pulled
across a rough horizontal surface by a force
applied to one of the blocks, as shown.
The acceleration of gravity is 9.8 m/s2
If each block has an acceleration of 2.5 m/s2
to the right, what is the magnitude of the
applied force?
Answer in units of N.
1
Expert's answer
2010-10-16T17:39:22-0400
There are the force of friction (μ* m *g) and the tension of the string T acting on each block acts.
For the first block the equation of the forces would be:
F- T - μ* m *g = ma
for the second block:
T - μ * m *g = ma

F - 2 μ * m *g = 2 ma
F = 2m (a + μ*g).

So for the completion of this question we have to know the coefficient of friction and the mass of blocks.